# Hybrid Opto-Electrical Excitation of Spin-Transfer Torque Nano-Oscillators for Advanced Computing

### v1

@@ III.2 C: Thermoelectric Effects in STNOs with Elliptical Nanopillars
STNOs with elliptical cross-sections exhibit in-plane magnetic anisotropy, stabilizing two energetically favorable antiparallel magnetization states along the major axis, thus enabling binary encoding (0 and 1). TMS measurements for two STNOs with dimensions of (eccentricity, e=0.96) and (e=0.87) are presented in Fig. 5. The smaller elliptical nanopillar () exhibits greater shape anisotropy and a stronger demagnetizing field, resulting in a higher energy barrier for magnetization switching. Consequently, a stronger external magnetic field is required to overcome this barrier, as evidenced by the increased switching field compared to the larger one (see Fig. 5). Increasing laser power reduces the switching field by lowering the energy barrier and decreasing the saturation magnetization, thereby weakening the demagnetizing field. Additionally, these elliptical STNOs, despite their smaller size relative to the circular 300 nm devices, exhibit a larger voltage difference (), indicating a higher contrast between the parallel (P) and antiparallel (AP) states.
@@
as a function of the magnetic field for varying laser powers and bias currents is shown in Fig. 6 for elliptical STNOs with different eccentricities (e). The STNO, with a smaller e, demonstrates lower magnetic field switching thresholds and higher thermovoltage output compared to devices with greater eccentricities. Interestingly, the measured displays small jumps, reminiscent of features commonly associated with vortex states. This suggests that elliptical nanopillars with smaller eccentricities may support vortex formation under specific laser power, potentially influencing device performance and stability. Our results demonstrate that arrays of STNOs with identical layer configurations but diverse cross-sectional geometries and dimensions on a single chip can generate a wide range of responses, which is critical for advanced computing applications.
@@ IV Conclusion
In summary, we have fabricated and optimized an array of spin-transfer torque nano-oscillators (STNOs) with both circular and elliptical cross-sections. A key aspect of this work is the implementation of a hybrid excitation scheme, enabling independent or simultaneous activation via AC laser illumination and DC bias current. Laser illumination alone induced substantial millivolt-range thermoelectric voltages via the tunnel magneto-Seebeck (TMS) effect, demonstrating CMOS compatibility. This breakthrough paves the way for simplified, energy-efficient neuromorphic chip architectures with reduced wiring complexity. When combined with a bias current, the bias-enhanced TMS (bTMS) effect was observed, characterized by giant thermovoltage outputs and novel two distinct phenomena: thermovoltage spikes and double-switching behavior. The thermovoltage spikes, corresponding to vortex state transitions, reveal exceptional sensitivity to magnetization changes and provide crucial insights into vortex core dynamics, including gyrotropic motion and pinning site jumps—fundamental aspects for STNO behavior. In contrast, the double-switching behavior likely originates from either instabilities in the fixed layer, leading to partial magnetization reversal, or highly nonlinear magnetization dynamics within the free layer.
Beyond their fundamental significance, the pulsed nature of the output voltage in both TMS and bTMS regimes resembles the spiking behavior of biological neurons, suggesting potential applications in systems inspired by neuromorphic computing. This behavior enables the integration of large arrays of nanopillars, each optically excited at distinct frequencies and sharing a common electrical output, creating an efficient frequency-multiplexing scheme. By employing frequency-selective devices in subsequent processing layers, this approach could reduce hardware complexity and enhance scalability Leroux et al. (2021). Furthermore, the combination of optical excitation and frequency-encoded spintronic signal processing offers opportunities for hybrid photonic-spintronic architectures, which may contribute to advancements in adaptive computing and signal processing. Future time-domain studies will be crucial to unravel the transient dynamics of pulsed thermovoltage and double-switching behavior, providing further insights into their underlying mechanisms and potential applications in neuromorphic systems, high-resolution sensing, and multistate memory storage. Leveraging their unique properties, hybrid laser-STNO systems hold great promise for bridging the gap between fundamental spintronic physics and next-generation computing technologies.

### v3

@@ III.2 Thermoelectric Effects in STNOs with Elliptical Nanopillars
STNOs with elliptical cross-sections exhibit in-plane magnetic anisotropy, stabilizing two energetically favorable antiparallel magnetization states along the major axis. This intrinsic anisotropy enables binary encoding (0 and 1), making these structures highly relevant for spintronic memory and logic applications. TMS measurements for two STNOs with dimensions of (eccentricity e = 0.96) and (e = 0.87) are presented in Fig. 5. The smaller elliptical nanopillar () exhibits greater shape anisotropy and a stronger demagnetizing field, resulting in a higher energy barrier for magnetization switching. Consequently, a stronger external magnetic field is required to overcome this barrier, as evidenced by the increased switching field compared to the larger one (see Fig. 5). Increasing laser power reduces the switching field by lowering the energy barrier and decreasing the saturation magnetization, thereby weakening the demagnetizing field. Additionally, these elliptical STNOs, despite their smaller size relative to the circular 300 nm devices, exhibit a larger voltage difference (), indicating a higher contrast between the parallel (P) and antiparallel (AP) states.
@@
as a function of the magnetic field for varying laser powers and bias currents is shown in Fig. 6 for elliptical STNOs with different eccentricities. The STNO, with a smaller eccentricity, demonstrates lower magnetic field switching thresholds and higher thermovoltage output compared to devices with greater eccentricities. Interestingly, the measured displays small jumps, reminiscent of features commonly associated with vortex states. This suggests that elliptical nanopillars with smaller eccentricities may support vortex formation under specific laser power, potentially influencing device performance and stability. Our results demonstrate that arrays of STNOs with identical layer configurations but diverse cross-sectional geometries and dimensions on a single chip can generate a wide range of responses, which is critical for advanced computing applications.
@@ IV Conclusion
In summary, we have fabricated and optimized an array of STNOs with both circular and elliptical cross-sections. A key aspect of this work is the implementation of a hybrid excitation scheme, enabling independent or simultaneous activation via AC laser illumination and DC bias current. Laser illumination alone induced substantial millivolt-range thermoelectric voltages via the TMS effect, demonstrating CMOS compatibility. This breakthrough paves the way for simplified, energy-efficient neuromorphic chip architectures with reduced wiring complexity. Incorporating a bias current further enhances the system’s response, leading to a bias-enhanced TMS (bTMS) effect with significantly increased thermoelectric outputs. The observed thermovoltage spikes provide direct insight into vortex state transitions, revealing exceptional sensitivity to magnetization changes, vortex core motion, and pinning site interactions—critical phenomena in STNO dynamics. Additionally, the double-switching behavior observed under specific biasing and laser power conditions suggests the presence of nonlinear magnetization dynamics or instabilities in the fixed layer, requiring further investigation to fully elucidate its origin.
Beyond their fundamental significance, the pulsed nature of the output voltage in both TMS and bTMS regimes resembles the spiking behavior of biological neurons, suggesting potential applications in systems inspired by neuromorphic computing. This behavior enables the integration of large arrays of nanopillars, each optically excited at distinct frequencies and sharing a common electrical output, creating an efficient frequency-multiplexing scheme. By employing frequency-selective devices in subsequent processing layers, this approach could reduce hardware complexity and enhance scalability Leroux et al. (2021). Furthermore, the combination of optical excitation and frequency-encoded spintronic signal processing offers opportunities for hybrid photonic-spintronic architectures, which may contribute to advancements in adaptive computing and signal processing. Future time-domain studies will be crucial to unravel the transient dynamics of pulsed thermovoltage and double-switching behavior, providing further insights into their underlying mechanisms and potential applications in neuromorphic systems, high-resolution sensing, and multistate memory storage. By leveraging their unique thermal, magnetic, and nonlinear dynamical properties, hybrid laser-STNO systems hold significant promise for bridging the gap between fundamental spintronic physics and next-generation computing technologies.